bestatcheer18:

The arithmetic mean (A) of two numbers (a and b) is given by the formula , and their geometric mean (G) is given by . Their harmonic mean (H) is given by the formula . Which formula correctly gives H in terms of a and b? A. H= 2ab/a+b B. H= square root of 2ab/a+b C. H= a+b/2ab D. H= ab/2(a+b)

Mercury:

\[A = \frac{ a+b }{ 2 }\] \[G = \sqrt{ab}\] \[G = \sqrt{AH}\] since you have two formulas for G, you can set sqrt(ab) and sqrt(AH) equal to each other \[\sqrt{ab}= \sqrt{AH}\] you can square both sides to get rid of the square root ab = AH now, you also know capital A in terms of lowercase a and b. plug in \(A = \frac{ a+b }{ 2 }\) into the equation, and solve for H.